### v1

# Omnidirectional Multi-Object Tracking

Abstract
Panoramic imagery, with its 360° field of view, offers comprehensive information to support Multi-Object Tracking (MOT) in capturing spatial and temporal relationships of surrounding objects. However, most MOT algorithms are tailored for pinhole images with limited views, impairing their effectiveness in panoramic settings. Additionally, panoramic image distortions, such as resolution loss, geometric deformation, and uneven lighting, hinder direct adaptation of existing MOT methods, leading to significant performance degradation. To address these challenges, we propose OmniTrack, an omnidirectional MOT framework that incorporates Tracklet Management to introduce temporal cues, FlexiTrack Instances for object localization and association, and the CircularStatE Module to alleviate image and geometric distortions. This integration enables tracking in large field-of-view scenarios, even under rapid sensor motion. To mitigate the lack of panoramic MOT datasets, we introduce the QuadTrack dataset—a comprehensive panoramic dataset collected by a quadruped robot, featuring diverse challenges such as wide fields of view, intense motion, and complex environments. Extensive experiments on the public JRDB dataset and the newly introduced QuadTrack benchmark demonstrate the state-of-the-art performance of the proposed framework. OmniTrack achieves a HOTA score of 26.92% on JRDB, representing an improvement of 3.43%, and further achieves 23.45% on QuadTrack, surpassing the baseline by 6.81%. The dataset and code will be made publicly available at https://github.com/xifen523/OmniTrack.
1 Introduction
Panoramic cameras, with a 360° Field of View (FoV), capture comprehensive surrounding information, making them essential for applications like autonomous driving [70, 10], robotic navigation [67, 63], and human-computer interaction [72, 29]. For small-scale mobile robots, such as quadrupedal robots, panoramic cameras are especially advantageous, allowing complete environmental awareness within a single compact setup, as illustrated in Fig. LABEL:fig:1(a).
Despite progress in Multi-Object Tracking (MOT), panoramic MOT remains underexplored. Existing MOT algorithms [14, 50], developed for pinhole cameras, struggle in panoramic settings due to inherent challenges like resolution loss, geometric distortion, and uneven color and brightness distribution when unfolded (Fig. LABEL:fig:1(d)). These challenges often lead to performance degradation when applying pinhole-based algorithms to panoramic images, limiting their effectiveness for panoramic scene perception.
To address these challenges, developing an MOT algorithm capable of comprehensive perception in panoramic images with large FoV is a pressing problem. To this end, this paper, for the first time, proposes an omnidirectional multi-object tracking framework, OmniTrack, specifically designed for such tasks in 360° panoramic imagery. OmniTrack unifies two mainstream MOT paradigms—Tracking-By-Detection (TBD) and End-To-End (E2E) tracking—and introduces a feedback mechanism that effectively reduces uncertainty in large FoV with rapid sensor motion, enabling fast and accurate target localization and association.
This framework consists of three core components: a CircularStatE Module, FlexiTrack Instance, and Tracklets Management. The CircularStatE Module is designed to mitigate wide-angle distortion and enhance consistency in lighting and color. The FlexiTrack Instance exploits the temporal continuity of objects, guiding the perception module to focus on key areas within the wide FoV and aiding in localization and association. This approach helps mitigate the difficulty of object localization in extra-wide FoV. The Tracklets Management module collects and manages trajectory data, providing prior knowledge to the FlexiTrack Instance. Through these components, OmniTrack unifies the two MOT paradigms: disabling data association within Tracklets Management results in an End-to-End tracker, OmniTrackE2E, while enabling association yields a TBD-style tracker, OmniTrackDA. By employing the same data association strategy, as shown in Fig. LABEL:fig:1 (c), the framework of OmniTrackDA achieves significantly stronger performance. Disabling both the FlexiTrack Instance and Tracklets Management reduces the system to a panoramic object detector, OmniTrackDet, as shown in Fig. LABEL:fig:1 (b).
Moreover, to support panoramic MOT research, we developed QuadTrack, a dataset collected with a $360^{\circ}{\times}70^{\circ}$ panoramic camera mounted on a quadrupedal robot. This mobile platform’s biomimetic gait introduces realistic, complex motion characteristics, challenging existing MOT algorithms. Collected across five campuses in two cities, QuadTrack includes $19,200$ images, encompassing a wide variety of dynamic, real-world scenarios. In contrast to typical MOT datasets [55, 8, 18, 5, 77, 15] that use static or linearly moving platforms, QuadTrack provides a new benchmark for evaluating MOT performance in large-FoV scenarios with rapid and non-linear sensor motion.
At a glance, our work makes the following contributions:
  • To address the gap in omnidirectional multi-object tracking, we propose OmniTrack, a novel framework that unifies both E2E and TBD tracking paradigms. This approach reduces uncertainty and enhances perceptual and association performance in large-FoV scenarios.
  • We present QuadTrack, a new panoramic MOT dataset with complex motion dynamics, providing a challenging benchmark for wide-FoV multi-object tracking.
  • Extensive experiments on JRDB and QuadTrack datasets show OmniTrack’s superior performance, achieving a $26.92\%$ HOTA on JRDB and $23.45\%$ on QuadTrack test splits, advancing the state-of-the-art in panoramic MOT.
2 Related Work
Panoramic scene understanding. Panoramic perception enables a holistic understanding of a 360° scene in a single shot [26, 13, 20, 23, 40, 39, 3]. Main areas include panoramic scene segmentation [65, 84, 10, 85, 74, 35, 36], panoramic estimation [4, 2, 68, 61, 12], panoramic layout estimation [78, 62, 48], panoramic generation [86, 69, 44], and panoramic flow estimation [63, 46], etc. [56, 41, 24, 29]. Researchers typically unfold panoramas into equirectangular projections or polyhedral projections to adapt algorithms designed for limited-FoV data [37, 68, 46]. They also apply techniques such as deformable convolutions to handle severe distortions in high-latitude regions [63, 80].
Recently, researchers have recognized the advantages of omnidirectional images for tracking, particularly their ability to maintain continuous observation of targets without the out-of-view issues present in limited field-of-view setups. Jiang et al. [38] propose a $500$FPS omnidirectional tracking system using a three-axis active vision mechanism to capture fast-moving objects in complex environments. The 360VOT benchmark [33] is introduced for omnidirectional object tracking, focusing on spherical distortions and object localization challenges. Huang et al. [34] present 360Loc for omnidirectional localization that tackles cross-device challenges by generating lower-FoV query frames from 360° data. Another work by Xu et al. [73] introduces an extended bounding FoV (eBFoV) representation to alleviate spherical distortions in panoramic videos. Unlike previous methods, this work first explores extremely challenging large-FoV and intense-motion panoramic tracking for mobile robots, e.g., aiming to enhance the robot’s spatiotemporal understanding of objects in its surroundings.
Multi-object tracking. Object tracking primarily follows two paradigms: Tracking-By-Detection (TBD) [14, 22, 58, 83, 32, 50, 45, 57] and End-To-End (E2E) [19, 47, 25, 79]. Among these, TBD is currently one of the most prevalent, with frameworks following the design principles of SORT [71]. First, the detection network [27, 11] is used to locate bounding boxes for objects, then the target’s current position is predicted based on its historical trajectory, and the predicted results are associated with detection results [43]. Many subsequent works have refined this approach: DeepSORT [46] introduced a ReID model to incorporate appearance information for association, and ByteTrack [81] designed a confidence-based, stage-wise association strategy. Other methods [1, 76, 21] introduced motion compensation modules to mitigate camera motion, and OC-SORT [9] optimized the motion estimation module. Additionally, E2E methods have continued to evolve. TrackFormer [53] and MOTR [79] proposed transformer-based, End-to-End tracking approaches. Recent improvements [82, 50] have enhanced detector performance and improved data association accuracy in occlusion scenarios. Unlike existing methods that focus on narrow-FoV pinhole camera data with linear sensor motion, we address the challenges of MOT in wide-FoV scenarios, tackling issues such as geometric distortion and complex motion.
Figure 1: The proposed OmniTrack pipeline. CSEM refers to the CircularStatE Module 3.4 , DA stands for data association, E2E denotes the End-to-End tracking paradigm, TBD refers to the Track-By-Detection tracking paradigm, Upd refers to updating tracks, Init to initializing tracks, and Del to deleting tracks.
1
Input: A Panoramic video/image sequence V
2
Output: Tracks $\mathcal{T}$ of the video/image sequence
3
4Initialization: $\mathcal{T}\leftarrow\emptyset$;
5
6Define the Initialize threshold $\mathcal{\tau}_{\mathcal{I}}$ ;
7 Define the Update threshold $\mathcal{\tau}_{\mathcal{U}}$ ;
8
9for frame $f_{k}$ in V do
       /* As shown in Fig. 1 */
10
11      $\{\mathcal{S}_{3},\mathcal{S}_{4},\mathcal{S}_{5}\}\leftarrow\texttt{Backbone}(f_{k})$ ;
12
13      $\mathcal{I}_{L}\leftarrow\texttt{CSEM}(\{\mathcal{S}_{3},\mathcal{S}_{4},\mathcal{S}_{5}\})$ ;
14
15      $\mathcal{I}_{F}\leftarrow\mathcal{T}_{f_{k-1}}$ ;
16
17      ${\color[rgb]{0.0,0.6,0.0}\mathcal{D}_{k}^{F}},\mathcal{D}_{k}^{L}\leftarrow\texttt{Decoder}({\color[rgb]{0.0,0.6,0.0}\mathcal{I}_{F}},\mathcal{I}_{L})$ ;
18
19      if DA then
             /* Data Association */
20
[…]
22
[…]
24
[…]
26
27      else
             /* End-to-End */
28
29            for $d$ in $\{\mathcal{D}_{k}^{F}\cup\mathcal{D}_{k}^{L}$}  do
30                   if $d\in\mathcal{D}_{k}^{F}\And d.score>\mathcal{\tau}_{\mathcal{U}}$ then
31                         $\texttt{Update}\leftarrow d$ ;
32
33                  if $d\in\mathcal{D}_{k}^{L}\And d.score>\mathcal{\tau}_{\mathcal{I}}$ then
34                         Initialize $\texttt{Initialize }\leftarrow d$ ;
35
36                  else
37                         $\texttt{Delete}\leftarrow d$ ;
38
39
[…]
41
42
43Return: $\mathcal{T}$
Algorithm 1 OmniTrack Inference Process
3 OmniTrack: Proposed Framework
In this section, we introduce OmniTrack, a panoramic multi-object tracking framework that addresses the unique challenges in wide-FoV images, including extensive search spaces, geometric distortion, resolution loss, and lighting inconsistencies. OmniTrack is designed with a feedback mechanism to iteratively refine object detection, integrating trajectory information back into the detector to enhance tracking stability across large-FoV scenes (Sec. 3.1). Specifically, to address these challenges, we propose three key components:
  • Tracklets Management (Sec. 3.2): Manages object trajectory lifecycles and provides temporal priors to the perception module.
  • FlexiTrack Instance (Sec. 3.3): Rapidly locates and associates objects across the panoramic view by leveraging temporal context.
  • CircularStatE Module (Sec. 3.4): Mitigates geometric distortion and improves consistency across the wide FoV, enhancing feature reliability.
3.1 Feedback Mechanism
The OmniTrack framework, illustrated in Fig. 1, differs from conventional MOT methods [81, 9, 21, 1], which separate object detection and data association. OmniTrack introduces a feedback mechanism that iteratively refines detections by integrating trajectory information back into the detector, enhancing stability for large-FoV tracking.
In traditional MOT, detection and association are decoupled, leading to higher entropy as each frame’s detection $H(x_{t})$ is calculated independently:
$H(x_{t})=-\sum_{i=1}^{n}P(x_{t}^{i})\log P(x_{t}^{i}),$ (1)
where $x_{t}^{i}$ denotes the position of the $i$-th target in frame $t$, with probability distribution $P(x_{t}^{i})$. The global association entropy $H(\{y_{t}\})$ depends on the joint probability distribution of target positions across all frames:
$H(\{y_{t}\})=-\sum_{i=1}^{n}$ $P(\{x_{1}^{i},x_{2}^{i},\dots,x_{T}^{i}\})$
$\times$ $\log P(\{x_{1}^{i},x_{2}^{i},\dots,x_{T}^{i}\}).$ (2)
The cumulative entropy across all frames, accounting for independent matching, is formulated as:
$H_{\text{independent}}=\sum_{t=1}^{T}H(x_{t})+H(\{y_{t}\}).$ (3)
In contrast, OmniTrack’s feedback mechanism allows detections from frame $t{-}1$ to inform those in frame $t$, reducing per-frame uncertainty. Specifically, the conditional entropy of frame $t$, given prior feedback $y_{t-1}$, is:
$H(x_{t}|y_{t-1})=-\sum_{i=1}^{n}P(x_{t}^{i}|y_{t-1}^{i})\log P(x_{t}^{i}|y_{t-1}^{i}).$ (4)
The total entropy with feedback becomes:
$H_{\text{feedback}}=\sum_{t=1}^{T}H(x_{t}|y_{t-1}),$ (5)
where $H_{\text{feedback}}{<}H_{\text{independent}}$, indicating a reduction in uncertainty over time. This feedback-driven approach therefore enhances tracking stability in large-FoV scenarios.
Figure 2: The proposed CircularStatE Module fuses multi-scale features to generate learnable instances. The DynamicSSM Block mitigates distortions in large-FoV images, enhancing feature stability across uneven lighting and color distributions.
3.2 Tracklets Management
To reduce uncertainty in target localization and association while incorporating temporal information, OmniTrack incorporates a Tracklets Management module. During training, this module caches temporal data for instances with confidence scores exceeding a threshold $\tau$, providing historical context to improve detection consistency in subsequent frames. During inference, Tracklets Management oversees trajectory lifecycle management by updating, deleting, or initializing instances based on their confidence scores. In scenarios without data association, trajectories are managed directly, forming OmniTrackE2E (Alg. 1, Lines 14-21). When data association is enabled, Tracklets Management utilizes TBD-based methods [9, 75] to enhance tracking, referred to as OmniTrackDA (Alg. 1, Lines 10-12)
3.3 FlexiTrack Instance
As described in Equ. (2), the global association entropy is significantly high under wide-FoV conditions, making the association task challenging. Inspired by human tracking behavior, which focuses on probable regions based on past movement, we estimate the target’s current position by leveraging its historical trajectory. This approach avoids global search across the entire field of view, which is particularly beneficial in large-scale perception tasks. Based on this insight, we introduce FlexiTrack Instance.
Each FlexiTrack Instance shares the Decoder network structure with Learnable Instances, consisting of a feature vector $\mathcal{X}{\in}\mathbb{R}^{128}$ and an anchor $\mathcal{Y}{\in}\mathbb{R}^{128}$, as shown in Fig. 1. By sharing the decoder, FlexiTrack Instances can seamlessly adapt to various MOT paradigms, enhancing flexibility and allowing integration across different approaches without additional modifications. To enhance robustness, noise is added to both feature vectors and anchors during training, minimizing dependency on historical data and improving generalization:
$\mathcal{X}^{\prime}=\mathcal{X}+\mathcal{N}_{X},\quad\mathcal{Y}^{\prime}=\mathcal{Y}+\mathcal{N}_{Y},$ (6)
where $\mathcal{N}_{X}$ and $\mathcal{N}_{Y}$ represent the noise components added to the feature vector and anchor, respectively. To initialize all FlexiTrack Instance, let $\mathcal{I_{F}}$ denote the set of initial instance, and $N$ the total number of trajectories. Each instance $\mathcal{I_{F}}$ is composed of a feature vector $\mathcal{X}_{i}$ and an anchor $\mathcal{Y}_{i}$, as:
=$\mathcal{I_{F}}=\left\{\mathcal{I_{F}}^{i}\mid\mathcal{I_{F}}^{i}=(\mathcal{X}^{\prime}_{i},\mathcal{Y}^{\prime}_{i}),i\in\{1,2,\dots,N\}\right\}.$ (7)
$\mathcal{X}^{\prime}_{i}{\in}\mathbb{R}^{d_{\mathcal{X}}}$ and $\mathcal{Y}^{\prime}_{i}{\in}\mathbb{R}^{d_{\mathcal{Y}}}$ are the feature vector and anchor of the $i$-th trajectory, with $d_{\mathcal{X}}{=}d_{\mathcal{Y}}{=}128$ representing their respective dimensions. This enables $\mathcal{I_{F}}$ to inherit trajectory information, guiding the perception module to quickly locate the object and establish temporal associations.
Data Domain
Datasets Cov. Pano. Platform Movement Trk Len No. Seq No. Smp No. T
KITTI MOT [28] n.a. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] n.a. 21 8k 749
Waymo [52] 220 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] 20s 103k 20m n.a.
nuScenes [8] 360 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] 20s 1000 40k n.a.
BDD100K MOT [77] n.a. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] 40s 2000 398k n.a.
SportsMOT [15] n.a. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] n.a. 240 150k 3401
DanceTrack [64] n.a. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] n.a. 100 105k 990
JRDB [51] 360 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] $\leq$117s 54 20k n.a.
MOT17 [55] n.a. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image] $\leq$85s 14 11k 1331
MOT20 [18] n.a. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] $\leq$133s 8 13k 3833
QuadTrack (ours) 360° [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] 60s 32 19k 332
Table 1: Typical datasets for 2D tracking. Abbreviations: [Uncaptioned image] (Autonomous Car), [Uncaptioned image] (Mobile Robot), [Uncaptioned image] (Quadruped Robot), [Uncaptioned image] (Internet images/videos), [Uncaptioned image] (Wheels), [Uncaptioned image] (Gait), [Uncaptioned image] (Stationary), Cov. (Coverage), Pano. (Panoramic image), Trk Len (Track Length), No. Seq (The number of sequences), No. Smp (The number of samples), and No. T (the number of tracks).
3.4 CircularStatE Module
The panoramic image provides an exceptionally wide FoV, capable of capturing 360° scenes. However, this inevitably introduces issues such as geometric distortions and inconsistencies in color and brightness in real-world high-dynamic-range scenes. To address these challenges, this paper proposes the CircularStatE Module, which alleviates distortions and improves the consistency of image features, thereby enhancing the performance of perception models.
The DynamicSSM Block, which is central to the CircularStatE Module, is responsible for mitigating distortions and refining the feature map. The operation is broken down into the following steps:
Distortion and Scale Calculation. The first step is to compute both the distortion and scale information from the input feature map $S_{4}$:
$\mathbf{d},\mathbf{s}=\mathcal{D}(S_{4}),\,\sigma(\mathcal{M}(S_{4})),$ (8)
where, $\mathbf{d}$ and $\mathbf{s}$ represent the distortion and scale, respectively, both of which have dimensions $\mathbb{R}^{B\times C\times W\times H}$.
Mitigate Distortion. To correct distortions, we apply a dynamic convolution $\mathcal{D}_{conv}$ to refine the feature map. The operation can be expressed as:
$\mathbf{D}=\mathcal{D}_{conv}(\mathbf{d}\odot\mathbf{s},S_{4}),$ (9)
where the symbol $\odot$ represents the Hadamard product, ensuring effective integration of scale adjustments.
Improve Consistency. Following distortion correction, a State Space Model (SSM) [17] is applied to enhance light and color consistency in the panoramic image. The input to this step is the output from the previous stage, denoted as $\mathbf{D}{\in}\mathbb{R}^{B\times C\times W\times H}$, and can be represented as follows:
${\mathbf{D^{\ast}}}[b,c,x,y]=\frac{1}{N}\sum_{d\in\{scan\}}F_{S6}(S_{d}(\mathbf{D}[b,c,x,y])),$ (10)
where $N$ represents the number of scans, $S_{d}$ represents the scanning function, and $F_{S6}$ is the transformation function for the S6 block [17].
Feature Fusion. Finally, the outputs from the dynamic convolution branch and the residual branch are fused. The fusion module $\mathcal{F}$ combines the refined feature map ${\mathbf{D^{\ast}}}$ with a processed version of $S_{4}$ (obtained via a CNN operation $\mathcal{C}(S_{4})$) to yield the final output feature map $\mathbf{F}$:
$\mathbf{F}=\mathcal{F}(\mathcal{C}(S_{4})\oplus{\mathbf{D^{\ast}}}).$ (11)
$\oplus$ denotes the feature fusion operation, combining details from both branches for optimal feature representation.
4 QuadTrack: a Dynamic 360° MOT Dataset
Most existing MOT datasets [55, 18, 64] are captured using pinhole cameras, which are characterized by a narrow-FoV and linear sensor motion. However, when wide-FoV capture devices experience even slight movements, the entire scene can change drastically, posing significant challenges for object tracking. QuadTrack addresses this challenge by providing a benchmark specifically designed to test MOT algorithms under dynamic, non-linear motion conditions. It enables the evaluation of algorithm robustness in tracking objects with panoramic, non-uniform motion.
4.1 Dataset Collection and Challenges
To acquire a dataset with a wide FoV and complex motion dynamics, we utilized a quadruped robot dog as the data collection platform. This platform was selected for its biomimetic gait, which emulates the natural locomotion patterns of quadrupedal animals, introducing additional challenges for motion tracking due to its inherent complexity and variability. The robot measures $70cm{\times}31cm{\times}40cm$, with a maximum payload capacity of $7kg$. It can navigate vertical obstacles up to $15cm$ and inclines up to $30^{\circ}$, making it highly maneuverable in everyday environments. With $12$ joint motors, the robot replicates realistic walking motions at speeds up to $2.5m/s$. For sensing, we used a Panoramic Annular Lens (PAL) camera to capture wide-angle scenes with a FoV of $360^{\circ}{\times}70^{\circ}$. The camera has a pixel size of $3.45{\mu}m{\times}3.45{\mu}m$, a resolution of $5$ million effective pixels, and supports a maximum output of $2448{\times}2048$ pixels at $40.5$FPS. Mounted on the quadruped robot (see Fig. 3 (b)), the camera ensures an unobstructed, optimal view. Using this platform, we conducted outdoor data collection across five campuses in two cities, capturing data in diverse environments.
With the biomimetic gait of the quadruped robot, the collected panoramic images naturally exhibited characteristic shaking, particularly along the Y-axis (Fig. 3 (c)). Compared to the JRDB dataset [51], our QuadTrack dataset introduces more complex motion challenges. Additionally, the data faces challenges such as uneven exposure, color inconsistencies due to the wide FoV, and increased motion blur, as rapid relative displacement between moving objects and the background intensifies the blurring effect. More details can be found in the supplementary.
Figure 3: (a) shows the bounding box (bbox) size distribution for the training and validation sets, whereas (b) depicts the data collection platform and panoramic camera setup. (c) compares the Y-axis pixel positions of trajectories between the QuadTrack (Refer to caption) and JRDB [51] (Refer to caption) datasets, illustrating the significant vertical motion of the sensor in QuadTrack.
4.2 Data Distribution and Comparative Analysis
Unlike existing panoramic MOT datasets [55, 18, 28], which rely on pinhole cameras, QuadTrack, as shown in Tab. 1, is the first to be captured using a single $360^{\circ}$ panoramic camera. With a wider FoV ($360^{\circ}{\times}70^{\circ}$), QuadTrack significantly differs from traditional MOT datasets [55, 18]. In contrast to autonomous driving datasets [8, 77, 52], which often feature more predictable motion, QuadTrack incorporates complex, biologically inspired gait movements. Moreover, unlike internet-sourced datasets [64, 15], QuadTrack is designed to better reflect real-world application scenarios. While many existing datasets [54, 52, 8, 77] consist of short video sequences, QuadTrack emphasizes long-term tracking, with each video lasting $60$ seconds. To further challenge data association, we downsampled the dataset to $10$FPS, resulting in $600$ frames per sequence, spread across $32$ sequences. In total, QuadTrack includes $19,200$ frames and $189,876$ bounding boxes.
As illustrated in Fig. 3 (a), the distribution of both the training and test sets is consistent, ensuring a reliable and balanced evaluation of MOT methods. This similarity in the distribution between the sets reduces the potential for bias and allows for a more accurate comparison of model performance across varying conditions. The trajectories depicted in Fig. 3 (c) highlight the increased complexity of multi-object tracking under wide FoV conditions. Notably, the motion along the Y-axis is significantly more intense compared to JRDB, further complicating the tracking process.
5 Experiments
5.1 Experiment Setup
[…]
E2E TrackFormer [53] 19.16 0.95 19.66 17.79
MOTRv2 [82] 18.22 0.93 19.30 12.30
OmniTrackE2E (ours) 21.56 0.94 22.87 25.01
TBD SORT [7] 23.49 0.90 26.11 24.59
DeepSORT [71] 22.15 0.95 23.46 24.88
ByteTrack [81] 25.00 0.86 27.95 26.59
Bot-SORT [1] 22.90 0.91 24.27 23.08
OC-SORT [9] 25.04 0.84 27.89 25.64
HybirdSORT [75] 25.01 0.85 27.82 25.03
DiffMOT [50] 19.96 0.95 20.26 20.05
OmniTrackDA (ours) 26.92 0.84 30.26 26.60
Table 2: Comparison with state-of-the-art methods on the JRDB test set [51].
[…]
E2E TrackFormer [53] 19.62 0.97 17.75 3.16
MOTRv2 [82] 16.42 0.96 17.08 -0.06
OmniTrackE2E (ours) 19.87 0.98 19.47 -5.89
TBD SORT [7] 14.57 0.98 15.60 4.81
DeepSORT [71] 21.16 0.96 22.56 5.12
ByteTrack [81] 20.66 0.94 22.56 8.68
Bot-SORT [1] 15.77 0.99 15.65 5.92
OC-SORT [9] 20.83 0.94 22.60 7.65
HybirdSORT [75] 16.64 0.96 17.38 6.79
DiffMOT [50] 16.40 0.97 16.62 6.21
OmniTrackDA (ours) 23.45 0.94 26.41 9.68
Table 3: Comparison with state-of-the-art methods on the QuadTrack test set.
[…]
Track-By- Detection (TBD) SORT [7] YOLO11 [66] (baseline) 25.83 29.56 0.915 31.02 27.62 24.51
OmniTrackDet (ours) 26.34 (+0.51) 31.11 (+1.55) 0.907 (-0.008) 34.21 (+3.19) 30.52 (+2.90) 22.96 (-1.55)
OmniTrackDA (ours) 29.44 (+3.10) 33.27 (+2.16) 0.927(+0.020) 33.44 (-0.77) 35.16 (+4.64) 25.06 (+2.10)
ByteTrack [81] YOLO11 [66] (baseline) 27.85 32.20 0.896 34.46 31.49 25.15
OmniTrackDet(ours) 28.14 (+0.29) 32.97 (+0.77) 0.870 (-0.026) 37.36 (+2.90) 32.94 (+1.45) 24.29 (-0.86)
OmniTrackDA (ours) 29.58 (+1.44) 34.54 (+1.57) 0.859 (-0.011) 38.14 (+0.78) 34.71 (+1.77) 25.49 (+1.20)
OC-SORT [9] YOLO11 [66] (baseline) 29.26 33.69 0.874 34.22 31.81 27.48
OmniTrackDet (ours) 29.43 (+0.17) 34.11 (+0.42) 0.851 (-0.023) 38.72 (+4.50) 34.48 (+2.67) 25.39 (-2.09)
OmniTrackDA (ours) 30.65 (+1.22) 34.83 (+0.72) 0.838 (-0.013) 36.37 (-2.35) 35.58 (+1.10) 26.76 (+1.37)
HybridSORT [75] YOLO11 [66] (baseline) 29.71 34.16 0.877 34.71 31.70 28.39
OmniTrackDet (ours) 30.00 (+0.29) 34.09 (-0.07) 0.853 (-0.024) 32.32 (-2.39) 35.02 (+3.32) 26.09 (-2.30)
OmniTrackDA (ours) 31.05 (+1.05) 36.06 (+1.97) 0.850 (-0.003) 38.13 (+5.81) 35.08 (+0.06) 27.78 (+1.69)
E2E TrackFormer [53] n.a. 22.22 23.38 0.959 23.83 30.30 16.93
MOTR [79] n.a. 19.78 23.25 0.928 25.44 25.51 15.61
MOTRv2 [82] n.a. 24.68 25.49 0.911 17.05 26.83 22.97
OmniTrackE2E (ours) n.a. 25.12 27.42 0.925 34.99 33.35 19.17
Table 4: Results on the JRDB validation set [51]. The first four groups compare methods under the TBD paradigm, whereas the last group presents a comparison under the E2E paradigm. In the TBD paradigm, each method is evaluated under three detection methods: the baseline with YOLO11 [66] as the detector, the OmniTrackDet detector, and OmniTrackDA. The numbers represent the improvement relative to the previous line’s method.
[…]
- - 0.01 0.00 1.00 0.00
3.80 1.91 0.99 -0.01
24.32 26.20 0.925 29.25
25.12 27.42 0.93 34.99
Table 5: Analysis of FlexiTrack Instance: $\textit{I}_{dn}$ represents an instance generated using Ground Truth (GT), whereas $\textit{I}_{ft}$ refers to a FlexiTrack Instance.
Datasets.
We conduct experiments on two datasets: JRDB [51] and QuadTrack. JRDB is a panoramic dataset designed for crowded human environments, comprising $10$ training sequences, $7$ validation sequences, and $27$ test sequences. The panoramic images in this dataset are stitched using a wheeled mobile robot equipped with five pinhole cameras. It includes both outdoor and indoor scenes, characterized by significant occlusion and the presence of small objects. Additionally, some objects exhibit rapid relative motion to the robot, which presents substantial challenges for MOT algorithms. Detailed information regarding the QuadTrack dataset is elaborated in Sec. 4.
Metrics.
We utilize the CLEAR metrics [6], which include MOTA, False Positives (FP), False Negatives (FN), and others, in conjunction with IDF1 [60], OPSA [51], and HOTA [49] to provide a comprehensive evaluation of tracking performance across various dimensions. MOTA emphasizes detector performance, while IDF1 evaluates the tracker’s ability to maintain consistent identities. In contrast, HOTA integrates association accuracy, and localization accuracy, establishing itself as an increasingly significant metric for assessing tracking algorithms.
Implementation details.
To enable a fair comparison of various MOT algorithms, we retrained models on the JRDB dataset. For End-To-End (E2E) algorithms [82, 53, 79], we trained using the default parameters from the source code on JRDB. For the MOT algorithms [81, 9, 7, 75] based on the TBD paradigm, we selected the advanced YOLO11-X [66] as the baseline detector for training on JRDB. Additionally, OmniTrackDet was obtained by masking the Track Management module after training the OmniTrack model and saving the detection results. The AdamW optimizer [42] was used, with the learning rate set to $10^{-5}$. For additional experimental details, please refer to the supplementary.
5.2 Comparison with State of the Art
Tracking on JRDB test set.
In Tab. 2, we compare our OmniTrack with state-of-the-art methods on the JRDB test set. Firstly, our approach significantly outperforms existing algorithms across all tracking metrics, whether in comparison with End-to-End or TBD paradigms. Specifically, OmniTrack achieves an impressive HOTA of $21.56\%$ and an IDF1 of $22.87\%$ within the End-to-End framework, surpassing the current state-of-the-art method, MOTRv2 [82], by $3.34\%$ and $3.57\%$, respectively. Furthermore, in the TBD paradigm, even under the same detector conditions, OmniTrack outperforms the state-of-the-art HybridSORT [75] by $1.91\%$ in HOTA and $2.44\%$ in IDF1, demonstrating its superior performance.
Tracking on QuadTrack test set.
In Tab. 3, we present a comparison between OmniTrack and state-of-the-art methods on the QuadTrack test set. This dataset is particularly challenging, characterized by a wide FoV and rapid, non-linear sensor motion, which introduces significant complexities for traditional MOT algorithms. Despite these challenges, our method outperforms existing approaches, achieving the highest HOTA scores: $19.87\%$ for the E2E group and $23.45\%$ for the TBD group.
5.3 Paradigm Comparison
Baseline.
To further validate the advantages of OmniTrack, we conducted comparisons based on the TBD and E2E paradigms, as shown in Tab. 4. In the TBD paradigm, we evaluated several baseline tracking algorithms [7, 81, 9, 75]. Each tracking method was compared under three different detection setups: using YOLO11-X [66] as the baseline detector, OmniTrackDet as the detector (representing traditional TBD tracking where detection and tracking are independent), and OmniTrackDA with a feedback mechanism for TBD tracking. In the E2E paradigm, we used MOTR [79] as the baseline for comparison.
Result.
In the same tracking method, OmniTrackDet consistently outperforms YOLO11-X [66], showing an average improvement of $0.2\%$ in HOTA and $0.6\%$ in IDF1. This highlights the superior performance of OmniTrackDet over YOLO11-X [66]. Furthermore, when comparing OmniTrackDet to OmniTrackDA, the latter shows an average increase of $1.7\%$ in HOTA and $1.4\%$ in IDF1, demonstrating the effectiveness of the feedback mechanism (Sec. 3.1). In the E2E paradigm, OmniTrackE2E achieved the best result HOTA of $25.12\%$ and IDF1 of $27.42\%$.
5.4 Ablation Study
Analysis of the FlexiTrack instance.
Tab. 5 compares experiments with and without denoise instances and FlexiTrack instances during the training phase. Experiments ➀ and ➁ demonstrate that FlexiTrack Instances are crucial for achieving the tracking objective. In Experiment ➂, we observe that denoise instances, generated from Ground Truth (GT), significantly improve the HOTA score by providing stronger guidance. Experiments ➂ and ➃ further show that incorporating FlexiTrack instances after using denoise instances leads to a further improvement in the HOTA score.
[…]
- - - 23.296 25.496 0.93415
MLP MLP MLP 21.951 23.535 0.92151
Conv Conv Conv 23.565 25.814 0.90931
24.724 26.886 0.91934
24.426 26.016 0.92819
24.539 26.506 0.92776
25.120 27.423 0.92512
Table 6: Ablation study on the CircularStatE module. $S_{3}$, $S_{4}$, and $S_{5}$ represent multi-scale features extracted from the backbone [30]. MLP refers to fully connected layers, Conv to convolutional layers. The symbol $\checkmark$ indicates the use of DynamicSSM 3.4
Analysis of the CircularStatE module.
In Tab. 6, we evaluate the effectiveness of DynamicSSM in the CircularStatE, comparing it with other common designs such as Conv and MLP. The results from experiments ➁, ➂, and ➃ demonstrate a clear advantage for DynamicSSM. Experiments ➄, ➅, and ➆ further show that applying DynamicSSM to $S_{4}$ yields the best performance. This is because $S_{4}$ contains both semantic and texture information, which makes it easier for DynamicSSM to extract distortion and color details, thus aiding in calibration.
Analysis of the initialization and update thresholds.
In OmniTrackE2E, we analyzed the impact of the initial threshold and updated threshold on tracking performance. As shown in Fig. 4, both the initial threshold and updated threshold achieved HOTA scores exceeding $25\%$ within the range of $0.1$ to $0.7$. This demonstrates that OmniTrackE2E is robust to threshold variations, eliminating the need for fine-tuning to achieve optimal results.
Comparison of end-to-end model training.
Method #Params FLOPs MACs Training Time $\downarrow$
TrackFormer [53] 44.01M 335G 167G 108 hours
MOTR [79] 43.91M 1421G 709G 80 hours
MOTRv2 [82] 41.65M 1395G 696G 130 hours
OmniTrackE2E (ours) 63.13M 762G 369G 20 hours
Table 7: Comparison of parameters, FLOPs, MACs, and training time for various end-to-end models on the JRDB dataset [51].
In Tab. 7, we compare the number of parameters and training time of OmniTrackE2E with existing End-to-End methods. Our method trains over four times faster than other End-to-End methods using default parameters on the JRDB dataset. This is achieved by implementing identity association through FlexiTrack Instances, which significantly simplifies the model design of the association component and alleviates the challenges associated with E2E model training.
Figure 4: Effects of the trajectory initialization threshold and update threshold on the HOTA metric in OmniTrackE2E.
6 Conclusion
This paper presents OmniTrack, a multi-object tracking framework tailored for panoramic images, effectively addressing key challenges like geometric distortion, low resolution, and lighting inconsistencies. Central to OmniTrack is a feedback mechanism that reduces uncertainty in large-FoV tracking. The framework incorporates Tracklets Management for temporal stability, FlexiTrack Instance for rapid localization and association, and the CircularStatE Module to mitigate distortion and improve visual consistency. Additionally, we present QuadTrack, a cross-campus multi-object tracking dataset collected using a quadruped robot to support dynamic motion scenarios. This challenging dataset is designed to advance research in omnidirectional perception for robotics. Experiments verify that OmniTrack achieves state-of-the-art performance on public JRDB and the established QuadTrack datasets, demonstrating its effectiveness in handling panoramic tracking tasks.
Limitations. While OmniTrack demonstrates strong performance, our approach is currently limited to 2D panoramic tracking without 3D capabilities, restricting depth perception in complex scenes. Additionally, the method is centered around a mobile robotic platform. Future work could consider extending to 3D panoramic MOT or exploring human-robot collaborative perception to enhance situational awareness.

### v2

@@
 To address these challenges, we propose OmniTrack, an omnidirectional MOT framework that incorporates Tracklet Management to introduce temporal cues, FlexiTrack Instances for object localization and association, and the CircularStatE Module to alleviate image and geometric distortions.
-This integration enables tracking in large field-of-view scenarios, even under rapid sensor motion.
-To mitigate the lack of panoramic MOT datasets, we introduce the QuadTrack dataset—a comprehensive panoramic dataset collected by a quadruped robot, featuring diverse challenges such as wide fields of view, intense motion, and complex environments.
+This integration enables tracking in panoramic field-of-view scenarios, even under rapid sensor motion.
+To mitigate the lack of panoramic MOT datasets, we introduce the QuadTrack dataset—a comprehensive panoramic dataset collected by a quadruped robot, featuring diverse challenges such as panoramic fields of view, intense motion, and complex environments.
 Extensive experiments on the public JRDB dataset and the newly introduced QuadTrack benchmark demonstrate the state-of-the-art performance of the proposed framework.
 OmniTrack achieves a HOTA score of 26.92% on JRDB, representing an improvement of 3.43%, and further achieves 23.45% on QuadTrack, surpassing the baseline by 6.81%.
-The dataset and code will be made publicly available at https://github.com/xifen523/OmniTrack.
+The established dataset and source code are available at https://github.com/xifen523/OmniTrack.
 1 Introduction
@@
 These challenges often lead to performance degradation when applying pinhole-based algorithms to panoramic images, limiting their effectiveness for panoramic scene perception.
-To address these challenges, developing an MOT algorithm capable of comprehensive perception in panoramic images with large FoV is a pressing problem.
+To address these challenges, developing an MOT algorithm capable of comprehensive perception in panoramic images with panoramic FoV is a pressing problem.
 To this end, this paper, for the first time, proposes an omnidirectional multi-object tracking framework, OmniTrack, specifically designed for such tasks in 360° panoramic imagery.
-OmniTrack unifies two mainstream MOT paradigms—Tracking-By-Detection (TBD) and End-To-End (E2E) tracking—and introduces a feedback mechanism that effectively reduces uncertainty in large FoV with rapid sensor motion, enabling fast and accurate target localization and association.
+OmniTrack unifies two mainstream MOT paradigms—Tracking-By-Detection (TBD) and End-To-End (E2E) tracking—and introduces a feedback mechanism that effectively reduces uncertainty in panoramic FoV with rapid sensor motion, enabling fast and accurate target localization and association.
 This framework consists of three core components: a CircularStatE Module, FlexiTrack Instance, and Tracklets Management.
 The CircularStatE Module is designed to mitigate wide-angle distortion and enhance consistency in lighting and color.
-The FlexiTrack Instance exploits the temporal continuity of objects, guiding the perception module to focus on key areas within the wide FoV and aiding in localization and association.
-This approach helps mitigate the difficulty of object localization in extra-wide FoV.
+The FlexiTrack Instance exploits the temporal continuity of objects, guiding the perception module to focus on key areas within the panoramic FoV and aiding in localization and association.
+This approach helps mitigate the difficulty of object localization in panoramic FoV.
 The Tracklets Management module collects and manages trajectory data, providing prior knowledge to the FlexiTrack Instance.
@@
 Collected across five campuses in two cities, QuadTrack includes $19,200$ images, encompassing a wide variety of dynamic, real-world scenarios.
-In contrast to typical MOT datasets [55, 8, 18, 5, 77, 15] that use static or linearly moving platforms, QuadTrack provides a new benchmark for evaluating MOT performance in large-FoV scenarios with rapid and non-linear sensor motion.
+In contrast to typical MOT datasets [55, 8, 18, 5, 77, 15] that use static or linearly moving platforms, QuadTrack provides a new benchmark for evaluating MOT performance in panoramic-FoV scenarios with rapid and non-linear sensor motion.
 At a glance, our work makes the following contributions:
   • To address the gap in omnidirectional multi-object tracking, we propose OmniTrack, a novel framework that unifies both E2E and TBD tracking paradigms.
-This approach reduces uncertainty and enhances perceptual and association performance in large-FoV scenarios.
-  • We present QuadTrack, a new panoramic MOT dataset with complex motion dynamics, providing a challenging benchmark for wide-FoV multi-object tracking.
+This approach reduces uncertainty and enhances perceptual and association performance in panoramic-FoV scenarios.
+  • We present QuadTrack, a new panoramic MOT dataset with complex motion dynamics, providing a challenging benchmark for panoramic-FoV multi-object tracking.
   • Extensive experiments on JRDB and QuadTrack datasets show OmniTrack’s superior performance, achieving a $26.92\%$ HOTA on JRDB and $23.45\%$ on QuadTrack test splits, advancing the state-of-the-art in panoramic MOT.
@@
 [73] introduces an extended bounding FoV (eBFoV) representation to alleviate spherical distortions in panoramic videos.
-Unlike previous methods, this work first explores extremely challenging large-FoV and intense-motion panoramic tracking for mobile robots, e.g., aiming to enhance the robot’s spatiotemporal understanding of objects in its surroundings.
+Unlike previous methods, this work first explores extremely challenging panoramic-FoV and intense-motion panoramic tracking for mobile robots, e.g., aiming to enhance the robot’s spatiotemporal understanding of objects in its surroundings.
 Multi-object tracking.
@@
 Recent improvements [82, 50] have enhanced detector performance and improved data association accuracy in occlusion scenarios.
-Unlike existing methods that focus on narrow-FoV pinhole camera data with linear sensor motion, we address the challenges of MOT in wide-FoV scenarios, tackling issues such as geometric distortion and complex motion.
+Unlike existing methods that focus on narrow-FoV pinhole camera data with linear sensor motion, we address the challenges of MOT in panoramic-FoV scenarios, tackling issues such as geometric distortion and complex motion.
 Figure 1: The proposed OmniTrack pipeline.
@@
 3 OmniTrack: Proposed Framework
-In this section, we introduce OmniTrack, a panoramic multi-object tracking framework that addresses the unique challenges in wide-FoV images, including extensive search spaces, geometric distortion, resolution loss, and lighting inconsistencies.
-OmniTrack is designed with a feedback mechanism to iteratively refine object detection, integrating trajectory information back into the detector to enhance tracking stability across large-FoV scenes (Sec.
+In this section, we introduce OmniTrack, a panoramic multi-object tracking framework that addresses the unique challenges in panoramic-FoV images, including extensive search spaces, geometric distortion, resolution loss, and lighting inconsistencies.
+OmniTrack is designed with a feedback mechanism to iteratively refine object detection, integrating trajectory information back into the detector to enhance tracking stability across panoramic-FoV scenes (Sec.
 3.1).
-Specifically, to address these challenges, we propose three key components:
+Specifically, we propose the OmniTrack framework, which consists of three key components:
   • Tracklets Management (Sec. 3.2): Manages object trajectory lifecycles and provides temporal priors to the perception module.
   • FlexiTrack Instance (Sec. 3.3): Rapidly locates and associates objects across the panoramic view by leveraging temporal context.
-  • CircularStatE Module (Sec. 3.4): Mitigates geometric distortion and improves consistency across the wide FoV, enhancing feature reliability.
+  • CircularStatE Module (Sec. 3.4): Mitigates geometric distortion and improves consistency across the panoramic FoV, enhancing feature reliability.
 3.1 Feedback Mechanism
 The OmniTrack framework, illustrated in Fig.
-1, differs from conventional MOT methods [81, 9, 21, 1], which separate object detection and data association.
-OmniTrack introduces a feedback mechanism that iteratively refines detections by integrating trajectory information back into the detector, enhancing stability for large-FoV tracking.
-In traditional MOT, detection and association are decoupled, leading to higher entropy as each frame’s detection $H(x_{t})$ is calculated independently:
+1, incorporates a feedback mechanism that iteratively refines detections by integrating trajectory information back into the detector.
+This mechanism operates on the principle of reducing information entropy, thereby enhancing stability in Panoramic-FoV and improving MOT performance.
+In traditional MOT [81, 9, 21, 1], detection and association are decoupled, leading to higher entropy as each frame’s detection $H(x_{t})$ is calculated independently:
 $H(x_{t})=-\sum_{i=1}^{n}P(x_{t}^{i})\log P(x_{t}^{i}),$ (1)
@@
 $H_{\text{feedback}}=\sum_{t=1}^{T}H(x_{t}|y_{t-1}),$ (5)
-where $H_{\text{feedback}}{<}H_{\text{independent}}$, indicating a reduction in uncertainty over time. This feedback-driven approach therefore enhances tracking stability in large-FoV scenarios.
+where $H_{\text{feedback}}{<}H_{\text{independent}}$, indicating a reduction in uncertainty over time. This feedback-driven approach thus enhances tracking stability in panoramic-FoV scenarios.
 Figure 2: The proposed CircularStatE Module fuses multi-scale features to generate learnable instances.
-The DynamicSSM Block mitigates distortions in large-FoV images, enhancing feature stability across uneven lighting and color distributions.
+The DynamicSSM Block mitigates distortions in panoramic-FoV images, enhancing feature stability across uneven lighting and color distributions.
 3.2 Tracklets Management
@@
 3.3 FlexiTrack Instance
-As described in Equ.
-(2), the global association entropy is significantly high under wide-FoV conditions, making the association task challenging.
-Inspired by human tracking behavior, which focuses on probable regions based on past movement, we estimate the target’s current position by leveraging its historical trajectory.
-This approach avoids global search across the entire field of view, which is particularly beneficial in large-scale perception tasks.
+As described in Eq.
+(2), the global association entropy is significantly high under panoramic-FoV conditions, making the association task challenging.
+Benefiting from the Feedback Mechanism (Sec.
+3.1), which integrates trajectory information into the detector to reduce information entropy.
+This approach eliminates the need for global search across the entire field of view, making it especially effective for panoramic-scale perception tasks.
 Based on this insight, we introduce FlexiTrack Instance.
-Each FlexiTrack Instance shares the Decoder network structure with Learnable Instances, consisting of a feature vector $\mathcal{X}{\in}\mathbb{R}^{128}$ and an anchor $\mathcal{Y}{\in}\mathbb{R}^{128}$, as shown in Fig.
+Each FlexiTrack Instance (see Fig.
+1) shares the Decoder network structure with Learnable Instances, consisting of a feature vector $\mathcal{X}{\in}\mathbb{R}^{128}$ and an anchor $\mathcal{Y}{\in}\mathbb{R}^{128}$, as shown in Fig.
 1.
@@
 where $\mathcal{N}_{X}$ and $\mathcal{N}_{Y}$ represent the noise components added to the feature vector and anchor, respectively.
-To initialize all FlexiTrack Instance, let $\mathcal{I_{F}}$ denote the set of initial instance, and $N$ the total number of trajectories.
-Each instance $\mathcal{I_{F}}$ is composed of a feature vector $\mathcal{X}_{i}$ and an anchor $\mathcal{Y}_{i}$, as:
+To initialize all FlexiTrack Instances, let $\mathcal{I_{F}}$ denote the set of initial instances, and $N$ the total number of trajectories.
+Each instance $\mathcal{I_{F}}^{i}$ is composed of a feature vector $\mathcal{X}_{i}$ and an anchor $\mathcal{Y}_{i}$, as:
 =$\mathcal{I_{F}}=\left\{\mathcal{I_{F}}^{i}\mid\mathcal{I_{F}}^{i}=(\mathcal{X}^{\prime}_{i},\mathcal{Y}^{\prime}_{i}),i\in\{1,2,\dots,N\}\right\}.$ (7)
@@
 (Coverage), Pano.
-(Panoramic image), Trk Len (Track Length), No.
+(Panoramic camera), Trk Len (Track Length), No.
 Seq (The number of sequences), No.
@@
 3.4 CircularStatE Module
-The panoramic image provides an exceptionally wide FoV, capable of capturing 360° scenes.
+The panoramic image provides an exceptionally panoramic FoV, capable of capturing 360° scenes.
 However, this inevitably introduces issues such as geometric distortions and inconsistencies in color and brightness in real-world high-dynamic-range scenes.
@@
 Most existing MOT datasets [55, 18, 64] are captured using pinhole cameras, which are characterized by a narrow-FoV and linear sensor motion.
-However, when wide-FoV capture devices experience even slight movements, the entire scene can change drastically, posing significant challenges for object tracking.
+However, when panoramic-FoV capture devices experience even slight movements, the entire scene can change drastically, posing significant challenges for object tracking.
 QuadTrack addresses this challenge by providing a benchmark specifically designed to test MOT algorithms under dynamic, non-linear motion conditions.
-It enables the evaluation of algorithm robustness in tracking objects with panoramic, non-uniform motion.
+It enables evaluating algorithm robustness in tracking objects with panoramic, non-uniform motion.
 4.1 Dataset Collection and Challenges
-To acquire a dataset with a wide FoV and complex motion dynamics, we utilized a quadruped robot dog as the data collection platform.
+To acquire a dataset with a panoramic FoV and complex motion dynamics, we utilized a quadruped robot dog as the data collection platform.
 This platform was selected for its biomimetic gait, which emulates the natural locomotion patterns of quadrupedal animals, introducing additional challenges for motion tracking due to its inherent complexity and variability.
@@
 For sensing, we used a Panoramic Annular Lens (PAL) camera to capture wide-angle scenes with a FoV of $360^{\circ}{\times}70^{\circ}$.
-The camera has a pixel size of $3.45{\mu}m{\times}3.45{\mu}m$, a resolution of $5$ million effective pixels, and supports a maximum output of $2448{\times}2048$ pixels at $40.5$FPS.
+The camera has a pixel size of $3.45{\mu}m{\times}3.45{\mu}m$, a resolution of $5$ million effective pixels, and supports a maximum output of $2048{\times}2048$ pixels at $40.5$FPS.
 Mounted on the quadruped robot (see Fig.
 3 (b)), the camera ensures an unobstructed, optimal view.
-Using this platform, we conducted outdoor data collection across five campuses in two cities, capturing data in diverse environments.
+Using this platform, the outdoor data collection spans morning, noon, afternoon, and evening, in diverse unconstrained environments across five campuses in two cities.
 With the biomimetic gait of the quadruped robot, the collected panoramic images naturally exhibited characteristic shaking, particularly along the Y-axis (Fig.
-3 (c)).
+3 (c) and (d)).
 Compared to the JRDB dataset [51], our QuadTrack dataset introduces more complex motion challenges.
-Additionally, the data faces challenges such as uneven exposure, color inconsistencies due to the wide FoV, and increased motion blur, as rapid relative displacement between moving objects and the background intensifies the blurring effect.
+Additionally, the data faces challenges such as uneven exposure, color inconsistencies due to the panoramic FoV, and increased motion blur, as rapid relative displacement between moving objects and the background intensifies the blurring effect.
 More details can be found in the supplementary.
 Figure 3: (a) shows the bounding box (bbox) size distribution for the training and validation sets, whereas (b) depicts the data collection platform and panoramic camera setup.
-(c) compares the Y-axis pixel positions of trajectories between the QuadTrack (Refer to caption) and JRDB [51] (Refer to caption) datasets, illustrating the significant vertical motion of the sensor in QuadTrack.
+(c) and (d) compare the normalized Y-axis pixel positions of trajectories between the QuadTrack (Refer to caption) and JRDB [51] (Refer to caption) datasets, illustrating the significant vertical motion of the sensor in QuadTrack.
 4.2 Data Distribution and Comparative Analysis
@@
 1, is the first to be captured using a single $360^{\circ}$ panoramic camera.
-With a wider FoV ($360^{\circ}{\times}70^{\circ}$), QuadTrack significantly differs from traditional MOT datasets [55, 18].
+With a panoramic FoV ($360^{\circ}{\times}70^{\circ}$), QuadTrack significantly differs from traditional MOT datasets [55, 18].
 In contrast to autonomous driving datasets [8, 77, 52], which often feature more predictable motion, QuadTrack incorporates complex, biologically inspired gait movements.
@@
 The trajectories depicted in Fig.
-3 (c) highlight the increased complexity of multi-object tracking under wide FoV conditions.
-Notably, the motion along the Y-axis is significantly more intense compared to JRDB, further complicating the tracking process.
+3 (c) and (d) highlight the increased complexity of multi-object tracking under panoramic FoV conditions.
+Notably, the motion along the Y-axis is significantly more intense compared to JRDB [51], further increasing the difficulty of object detection and association.
 5 Experiments
@@
 OC-SORT [9] 25.04 0.84 27.89 25.64
-HybirdSORT [75] 25.01 0.85 27.82 25.03
+HybridSORT [75] 25.01 0.85 27.82 25.03
 DiffMOT [50] 19.96 0.95 20.26 20.05
@@
 OC-SORT [9] 20.83 0.94 22.60 7.65
-HybirdSORT [75] 16.64 0.96 17.38 6.79
+HybridSORT [75] 16.64 0.96 17.38 6.79
 DiffMOT [50] 16.40 0.97 16.62 6.21
@@
 […]
-Track-By- Detection (TBD) SORT [7] YOLO11 [66] (baseline) 25.83 29.56 0.915 31.02 27.62 24.51
-OmniTrackDet (ours) 26.34 (+0.51) 31.11 (+1.55) 0.907 (-0.008) 34.21 (+3.19) 30.52 (+2.90) 22.96 (-1.55)
-OmniTrackDA (ours) 29.44 (+3.10) 33.27 (+2.16) 0.927(+0.020) 33.44 (-0.77) 35.16 (+4.64) 25.06 (+2.10)
-ByteTrack [81] YOLO11 [66] (baseline) 27.85 32.20 0.896 34.46 31.49 25.15
-OmniTrackDet(ours) 28.14 (+0.29) 32.97 (+0.77) 0.870 (-0.026) 37.36 (+2.90) 32.94 (+1.45) 24.29 (-0.86)
-OmniTrackDA (ours) 29.58 (+1.44) 34.54 (+1.57) 0.859 (-0.011) 38.14 (+0.78) 34.71 (+1.77) 25.49 (+1.20)
-OC-SORT [9] YOLO11 [66] (baseline) 29.26 33.69 0.874 34.22 31.81 27.48
-OmniTrackDet (ours) 29.43 (+0.17) 34.11 (+0.42) 0.851 (-0.023) 38.72 (+4.50) 34.48 (+2.67) 25.39 (-2.09)
-OmniTrackDA (ours) 30.65 (+1.22) 34.83 (+0.72) 0.838 (-0.013) 36.37 (-2.35) 35.58 (+1.10) 26.76 (+1.37)
-HybridSORT [75] YOLO11 [66] (baseline) 29.71 34.16 0.877 34.71 31.70 28.39
-OmniTrackDet (ours) 30.00 (+0.29) 34.09 (-0.07) 0.853 (-0.024) 32.32 (-2.39) 35.02 (+3.32) 26.09 (-2.30)
-OmniTrackDA (ours) 31.05 (+1.05) 36.06 (+1.97) 0.850 (-0.003) 38.13 (+5.81) 35.08 (+0.06) 27.78 (+1.69)
-E2E TrackFormer [53] n.a. 22.22 23.38 0.959 23.83 30.30 16.93
-MOTR [79] n.a. 19.78 23.25 0.928 25.44 25.51 15.61
-MOTRv2 [82] n.a. 24.68 25.49 0.911 17.05 26.83 22.97
-OmniTrackE2E (ours) n.a. 25.12 27.42 0.925 34.99 33.35 19.17
+Track-By- Detection (TBD) SORT [7] YOLO11 [66] (baseline) 25.83 29.56 0.915 31.02 27.62 24.51 49.18
+OmniTrackDet (ours) 26.34 (+0.51) 31.11 (+1.55) 0.907 (-0.008) 34.21 (+3.19) 30.52 (+2.90) 22.96 (-1.55) 12.14
+OmniTrackDA (ours) 29.44 (+3.10) 33.27 (+2.16) 0.927(+0.020) 33.44 (-0.77) 35.16 (+4.64) 25.06 (+2.10) 11.78
+ByteTrack [81] YOLO11 [66] (baseline) 27.85 32.20 0.896 34.46 31.49 25.15 50.36
+OmniTrackDet(ours) 28.14 (+0.29) 32.97 (+0.77) 0.870 (-0.026) 37.36 (+2.90) 32.94 (+1.45) 24.29 (-0.86) 12.24
+OmniTrackDA (ours) 29.58 (+1.44) 34.54 (+1.57) 0.859 (-0.011) 38.14 (+0.78) 34.71 (+1.77) 25.49 (+1.20) 11.83
+OC-SORT [9] YOLO11 [66] (baseline) 29.26 33.69 0.874 34.22 31.81 27.48 46.33
+OmniTrackDet (ours) 29.43 (+0.17) 34.11 (+0.42) 0.851 (-0.023) 38.72 (+4.50) 34.48 (+2.67) 25.39 (-2.09) 11.59
+OmniTrackDA (ours) 30.65 (+1.22) 34.83 (+0.72) 0.838 (-0.013) 36.37 (-2.35) 35.58 (+1.10) 26.76 (+1.37) 11.13
+HybridSORT [75] YOLO11 [66] (baseline) 29.71 34.16 0.877 34.71 31.70 28.39 44.34
+OmniTrackDet (ours) 30.00 (+0.29) 34.09 (-0.07) 0.853 (-0.024) 32.32 (-2.39) 35.02 (+3.32) 26.09 (-2.30) 11.65
+OmniTrackDA (ours) 31.05 (+1.05) 36.06 (+1.97) 0.850 (-0.003) 38.13 (+5.81) 35.08 (+0.06) 27.78 (+1.69) 10.96
+E2E TrackFormer [53] n.a. 22.22 23.38 0.959 23.83 30.30 16.93 7.38
+MOTR [79] n.a. 19.78 23.25 0.928 25.44 25.51 15.61 12.73
+MOTRv2 [82] n.a. 24.68 25.49 0.911 17.05 26.83 22.97 13.01
+OmniTrackE2E (ours) n.a. 25.12 27.42 0.925 34.99 33.35 19.17 11.64
 Table 4: Results on the JRDB validation set [51].
@@
 The numbers represent the improvement relative to the previous line’s method.
+The FPS metric is measured on a single RTX 3090 GPU with an image resolution of $4160\times 480$.
 […]
@@
 3.80 1.91 0.99 -0.01
-24.32 26.20 0.925 29.25
+24.32 26.20 0.93 29.25
 25.12 27.42 0.93 34.99
@@
 Metrics.
-We utilize the CLEAR metrics [6], which include MOTA, False Positives (FP), False Negatives (FN), and others, in conjunction with IDF1 [60], OPSA [51], and HOTA [49] to provide a comprehensive evaluation of tracking performance across various dimensions.
-MOTA emphasizes detector performance, while IDF1 evaluates the tracker’s ability to maintain consistent identities.
-In contrast, HOTA integrates association accuracy, and localization accuracy, establishing itself as an increasingly significant metric for assessing tracking algorithms.
+We employ the CLEAR metrics [6], including MOTA, DetA, and AssA, alongside IDF1 [60], OPSA [51], and HOTA [49] for a comprehensive tracking performance evaluation.
+MOTA is primarily influenced by detector performance, IDF1 measures identity preservation, and HOTA integrates association and localization accuracy, making it increasingly pivotal for tracking assessment.
 Implementation details.
@@
 3, we present a comparison between OmniTrack and state-of-the-art methods on the QuadTrack test set.
-This dataset is particularly challenging, characterized by a wide FoV and rapid, non-linear sensor motion, which introduces significant complexities for traditional MOT algorithms.
+This dataset is particularly challenging, characterized by a panoramic FoV and rapid, non-linear sensor motion, which introduces significant complexities for traditional MOT algorithms.
 Despite these challenges, our method outperforms existing approaches, achieving the highest HOTA scores: $19.87\%$ for the E2E group and $23.45\%$ for the TBD group.
@@
 Result.
-In the same tracking method, OmniTrackDet consistently outperforms YOLO11-X [66], showing an average improvement of $0.2\%$ in HOTA and $0.6\%$ in IDF1.
-This highlights the superior performance of OmniTrackDet over YOLO11-X [66].
-Furthermore, when comparing OmniTrackDet to OmniTrackDA, the latter shows an average increase of $1.7\%$ in HOTA and $1.4\%$ in IDF1, demonstrating the effectiveness of the feedback mechanism (Sec.
-3.1).
+In the TBD method, OmniTrackDet consistently outperforms YOLO11-X [66], showing an average improvement of $0.2\%$ in HOTA and $0.6\%$ in IDF1.
+Despite OmniTrackDet not having a speed advantage, it achieves notable improvements in accuracy.
+Furthermore, when comparing OmniTrackDet to OmniTrackDA, the latter shows an average increase of $1.7\%$ in HOTA and $1.4\%$ in IDF1, demonstrating the effectiveness of the feedback mechanism.
 In the E2E paradigm, OmniTrackE2E achieved the best result HOTA of $25.12\%$ and IDF1 of $27.42\%$.
@@
 Experiments ➄, ➅, and ➆ further show that applying DynamicSSM to $S_{4}$ yields the best performance.
-This is because $S_{4}$ contains both semantic and texture information, which makes it easier for DynamicSSM to extract distortion and color details, thus aiding in calibration.
+where $S_{5}$, $S_{4}$, and $S_{3}$ impact MOT results.
+Since $S_{4}$ contains both high-level semantic and low-level geometric features, its effect is the most pronounced.
 Analysis of the initialization and update thresholds.
@@
 This paper presents OmniTrack, a multi-object tracking framework tailored for panoramic images, effectively addressing key challenges like geometric distortion, low resolution, and lighting inconsistencies.
-Central to OmniTrack is a feedback mechanism that reduces uncertainty in large-FoV tracking.
+Central to OmniTrack is a feedback mechanism that reduces uncertainty in panoramic-FoV tracking.
 The framework incorporates Tracklets Management for temporal stability, FlexiTrack Instance for rapid localization and association, and the CircularStatE Module to mitigate distortion and improve visual consistency.
@@
 Future work could consider extending to 3D panoramic MOT or exploring human-robot collaborative perception to enhance situational awareness.
+Acknowledgment
+This work was supported in part by the National Natural Science Foundation of China (No.
+62473139 and No.
+12174341), in part by Zhejiang Provincial Natural Science Foundation of China (Grant No.
+LZ24F050003), and in part by Shanghai SUPREMIND Technology Co.
+Ltd.